OpenStudy (anonymous):

Use the substitution method to solve the system of equations. 2x – 7y = 1 x – 8y = –4

OpenStudy (anonymous):

x = 8y - 4 2(8y - 4) - 7y = 1 16y - 8 - 7y = 1 9y = 9 y = 1 x = 8(1) - 4) x = 4 (4, 1)

OpenStudy (anonymous):

x-8y=-4 x=8y-4 substitute x in 1st equation to get y value then substitute y value in one of the equation

OpenStudy (anonymous):

to use substitution we need to solve one of those equations for one of the variables. lets solve the second equation for x since that looks easy. we get: x=8y-4 now we can plug that into our original equation

OpenStudy (anonymous):

by orginal i meant first. that wil give us: 2(8y-4)-7y=1 16y-8-7y=1 9y=9 y=1

OpenStudy (anonymous):

plug that back in to your solution for x and get: x=8-4 x=4 so your answer is x=4 y=1